OpenStudy (anonymous):

PLEASE HELP!!! f(x) = x^2 + 4x − 1 in vertex form

OpenStudy (anonymous):

@Nurali

OpenStudy (butterflydreamer):

hint: you want to use completing the square ^_^

OpenStudy (anonymous):

I don't know how to do that.

OpenStudy (butterflydreamer):

 So now you want to plug b = 4 into the formula: \[(\frac{ b }{ 2})^{2} \] \[= (\frac{ 4}{ 2 }) ^2 = 2^2 = 4 \] Therefore to complete the square we have to add 4 \[f(x) = x^2 + 4x - 1 \] \[= (x^2 + 4x + 4) - 1 \] \[= (x + 2)^2 -1\]